OpenStudy (anonymous):

check whether the given number is a solution of the equation or inequality: 9+4y=17;1

OpenStudy (texaschic101):

9 + 4y = 17 4y = 17 - 9 4y = 8 y = 2 does that answer your question

OpenStudy (texaschic101):

is there more to this question ?

OpenStudy (anonymous):

so the right one is 17 or 1

OpenStudy (anonymous):

wait I got it thanks

OpenStudy (texaschic101):

oh.....I misunderstood the question....hold on....wait....I do not understand your question. Are you asking what 9 + 4y equals

OpenStudy (anonymous):

yes

OpenStudy (texaschic101):

9 + 4y = 0 and you want to solve for y ??

OpenStudy (texaschic101):

or 9 + 4y = 17 and you want to solve for y ?

OpenStudy (anonymous):

Okay in the book it says check whether the given number is a solution of the equation

OpenStudy (anonymous):

so is 17;1 it or what

OpenStudy (texaschic101):

9 + 4y = 17 y = 2 The given number 1 is not a solution

OpenStudy (anonymous):

i also have another question like this one i am lost on it too

OpenStudy (texaschic101):

lets see it

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

\[2p-1\ge7;3\]

OpenStudy (texaschic101):

2p - 1 >= 7 2p >= 7 + 1 2p >= 8 p >= 4 3 is not a solution

OpenStudy (anonymous):

ok so i write down what you did?

OpenStudy (anonymous):

and is 3 the solution

OpenStudy (texaschic101):

sorry...3 is not a solution. If you want to show your work, you can write it down

OpenStudy (anonymous):

ok so what is the solution

OpenStudy (texaschic101):

the solution in this inequality is any number greater then 4......(5,6,7,8,9,etc)

OpenStudy (anonymous):

wait what you done in your work was less than and not less than or equal too

OpenStudy (anonymous):

2p - 1 >= 7

OpenStudy (texaschic101):

p is greater then or equal to 4......so count 4 also answers could be (4,5,6,7,8 etc)

OpenStudy (anonymous):

Ok so what should i put as my work :2p - 1 >= 7 2p >= 7 + 1 2p >= 8 p >= 4

OpenStudy (texaschic101):

yes

OpenStudy (anonymous):

OK THANKS!! i really hate algebra but have to take it

OpenStudy (texaschic101):

I understand......I hate geometry

OpenStudy (texaschic101):

do you have anymore ?

OpenStudy (anonymous):

nope not like that

OpenStudy (texaschic101):

ok.....good luck with algebra

OpenStudy (anonymous):

thanks i wish i had some one there to explain like you

OpenStudy (texaschic101):

well....I didn't do much explaining..do you want me too ?

OpenStudy (anonymous):

nope you did just fine

OpenStudy (texaschic101):

thanks....if you need more help, just tag me :)

OpenStudy (anonymous):

ok Thanks